I need assistance in setting this up and solving. Thanks
Elevation E, in meters, above sea leve at which the boiling point of water is
t degrees Celsius is given by the formula
E=1000(100-t) + 580(100-t)^2
At what elevation is the boing point 99.2 degrees Celsius?

again thank you!

Answer by ankor@dixie-net.com(22740)

Elevation E, in meters, above sea leve at which the boiling point of water is
t degrees Celsius is given by the formula:
:
E = 1000(100-t) + 580(100-t)^2
:
At what elevation is the boing point 99.2 degrees Celsius?
:
Substitute 99.2 for t in the formula
:
E = 1000(100 - 99.2) + 580(100 - 99.2)^2
:
E = 1000(.8) + 580(.8)^2
:
E = 800 + 580(.64)
:
E = 800 + 371.2
:
E = 1171.2 meters above sea level